# Omni-LIVO: Robust RGB-Colored Multi-Camera Visual-Inertial-LiDAR Odometry via Photometric Migration and ESIKF Fusion

### v2

@@ Abstract
Wide field-of-view (FoV) LiDAR sensors provide dense geometry across large environments, but existing LiDAR-inertial-visual odometry (LIVO) systems generally rely on a single camera, limiting their ability to fully exploit LiDAR-derived depth for photometric alignment and scene colorization. We present Omni-LIVO, a tightly coupled multi-camera LIVO system that leverages multi-view observations to comprehensively utilize LiDAR geometric information across extended spatial regions. Omni-LIVO introduces a Cross-View direct alignment strategy that maintains photometric consistency across non-overlapping views, and extends the Error-State Iterated Kalman Filter (ESIKF) with multi-view updates and adaptive covariance. The system is evaluated on public benchmarks and our custom dataset, showing improved accuracy and robustness over state-of-the-art LIVO, LIO, and visual-inertial SLAM baselines. Code and dataset will be released upon publication.
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) has become essential for autonomous systems operating in unknown environments. Traditional methods are categorized by sensor types: vision-based SLAM [1, 2] provides rich visual information but struggles in low-light or texture-poor scenes, while LiDAR-based SLAM [3, 4, 5] offers robust geometry but degrades in repetitive environments. IMUs, though environment-independent, suffer from long-term drift.
@@
Unlike stereo systems requiring overlapping views and feature matching, the key challenges are: maintaining photometric continuity across non-overlapping views through temporal patch migration as patches disappear from one camera and reappear in another, and adaptive covariance modeling for heterogeneous measurement reliability across different views [13, 14]. Directly leveraging LiDAR-derived depth for multi-camera sparse direct alignment maintains efficient performance while enabling accurate visual point selection and extended spatial coverage beyond single-camera constraints.
We present Omni-LIVO, a tightly coupled multi-camera LIVO system that extends FAST-LIVO2 with multi-view photometric constraints.
@@
    Complete open-source system: A tightly coupled multi-camera LIVO framework with comprehensive evaluation on public benchmarks and custom datasets.
@@ II-C Multi-Camera Visual-Inertial Systems
Multi-camera visual-inertial SLAM has gained attention for extended spatial coverage. MIMC-VINS [20] supports arbitrary camera numbers using multi-state constraint Kalman filtering. BAMF-SLAM [21] integrates multiple fisheye cameras with IMU through bundle adjustment. MAVIS [12] extends to partially overlapped multi-camera systems. These systems operate in the visual-inertial domain without leveraging LiDAR geometric structure.
@@ II-D Direct Visual Methods
Direct visual methods minimize photometric error without feature extraction. DTAM [22], LSD-SLAM [1], and DSO [23] established foundations for dense and sparse tracking. MultiCol-SLAM [24] extended direct methods to multi-camera systems. This approach suits multi-camera LIVO by avoiding feature correspondence complexity across non-overlapping views.
@@ III System Overview And Methodology
The Omni-LIVO system comprises four modules: Sensor Acquisition, Preprocessing (IMU integration and temporal alignment), ESIKF (fusing LiDAR and multi-view photometric constraints), and Unified Mapping (voxel representation). The multi-camera setup preserves photometric consistency via Cross-View temporal migration and adaptive ESIKF.
@@ III-B Multi-Sensor Data Alignment
Building upon FAST-LIVO2 [10], we extend temporal alignment to multi-camera configurations [25, 26]. The system forms temporally coherent measurement packets:
@@ III-C Cross-View Direct Method
$\mathbf{r}_{VIO}=\left[\begin{array}[]{c}\mathbf{r}^{1}_{\mathrm{direct}}\\ \vdots\\ \mathbf{r}^{N_{c}}_{\mathrm{direct}}\\ \mathbf{r}^{1\to 2}_{\mathrm{migration}}\\ \vdots\\ \mathbf{r}^{i\to j}_{\mathrm{migration}}\end{array}\right]$ (13)
@@ III-E Iterative Multi-Camera State Estimation with ESIKF
We extend the standard ESIKF to integrate multi-view photometric constraints with LiDAR measurements through sequential update and adaptive covariance, as shown in Fig. 5. The state is first updated with LiDAR point-to-plane residuals, then sequentially updated with multi-camera photometric residuals. Multi-camera residuals and Jacobians are stacked across all cameras and processed through iterative ESIKF updates until convergence.
@@
where $\bar{e}^{(t-1)}$ is the mean squared photometric error from the previous frame, $\sigma_{\min}$ and $\sigma_{\max}$ are the minimum and maximum covariance scaling bounds (set to $100.0$ and $1000.0$ respectively), and $e_{\max}$ is the maximum error threshold for scaling. Temporal smoothing ensures stability:
@@
Algorithm 1 summarizes the complete implementation. Algorithm-specific notation: $\mathbf{r}_{LIO}$, $\mathbf{H}_{LIO}$ are LiDAR-inertial residuals and Jacobians; $\delta\hat{\mathbf{x}}$ is state increment; $\ell_{\max}=10$ is maximum iterations. For multi-camera VIO: $\mathbf{r}_{c}^{(\ell)}$, $\mathbf{H}_{c}^{(\ell)}$ are per-camera residuals and Jacobians at iteration $\ell$; $\mathbf{r}^{(\ell)}_{VIO}$, $\mathbf{H}^{(\ell)}_{VIO}$ are stacked terms across all cameras; $\check{\mathbf{x}}^{(\ell)}$ is linearization point; $e^{(\ell)}$ is mean squared error at iteration $\ell$; $e_{\text{last}}$ stores the previous iteration’s error for convergence checking; $n_{\text{meas}}$ is total number of measurements; $\delta\mathbf{R}^{(\ell)}$, $\delta\mathbf{t}^{(\ell)}$ are rotation and translation increments.
@@
1: IMU propagation: $\hat{\mathbf{x}}^{-}$, $\mathbf{P}^{-}$; LiDAR plane extraction
2:LiDAR update: Iterative ESIKF with $\mathbf{r}_{LIO}$, $\mathbf{H}_{LIO}$
3: Select patches from cameras $c\in[1,N_{c}]$; Compute $\tilde{\sigma}^{(t)}$ via Eq. (17)-(18)
4:VIO update: $\ell\leftarrow 0$, $\delta\hat{\mathbf{x}}^{(0)}\leftarrow\mathbf{0}$, $e_{\text{last}}\leftarrow\infty$
5:repeat
6:  $\ell\leftarrow\ell+1$; Update $\check{\mathbf{x}}^{(\ell)}=\hat{\mathbf{x}}^{-}\boxplus\delta\hat{\mathbf{x}}^{(\ell-1)}$
7:  for camera $c=1,\ldots,N_{c}$ do
8:   Compute $\mathbf{r}_{c}^{(\ell)}$, $\mathbf{H}_{c}^{(\ell)}$ with cross-camera constraints
9:  end for
10:  Stack $\mathbf{r}^{(\ell)}_{VIO}=[\mathbf{r}_{1}^{(\ell)};\ldots;\mathbf{r}_{N_{c}}^{(\ell)}]$, $\mathbf{H}^{(\ell)}_{VIO}=[\mathbf{H}_{1}^{(\ell)};\ldots;\mathbf{H}_{N_{c}}^{(\ell)}]$; Compute $e^{(\ell)}=\|\mathbf{r}^{(\ell)}_{VIO}\|^{2}/n_{\text{meas}}$
11:  if $e^{(\ell)}\leq e_{\text{last}}$ then
12:   $e_{\text{last}}\leftarrow e^{(\ell)}$; $\mathbf{K}=((\mathbf{H}^{(\ell)}_{VIO})^{\mathrm{T}}\mathbf{H}^{(\ell)}_{VIO}+\mathbf{P}^{-1}/\tilde{\sigma}^{(t)})^{-1}$
13:   $\delta\hat{\mathbf{x}}^{(\ell)}=\mathbf{K}((\mathbf{H}^{(\ell)}_{VIO})^{\mathrm{T}}\mathbf{r}^{(\ell)}_{VIO}-\mathbf{P}^{-1}\delta\hat{\mathbf{x}}^{(\ell-1)}/\tilde{\sigma}^{(t)})$; $\hat{\mathbf{x}}^{-}\leftarrow\hat{\mathbf{x}}^{-}\boxplus\delta\hat{\mathbf{x}}^{(\ell)}$
14:  else
15:   break
16:  end if
17:until $\|\delta\mathbf{R}^{(\ell)}\|<0.001^{\circ}$ and $\|\delta\mathbf{t}^{(\ell)}\|<0.01$$m$ or $\ell\geq\ell_{\max}$
[…]
19:return $\hat{\mathbf{x}}_{k|k}$, $\mathbf{P}_{k|k}$
@@ IV Experimental Evaluation
We evaluate the proposed method on Hilti SLAM Challenge [29] and Newer College datasets [30], measuring absolute trajectory error (RMSE) against ground truth and comparing with state-of-the-art methods. We also collected custom dataset using LIVOX MID360 and four hardware-triggered cameras.
@@ IV-A Datasets and Baselines
Public Benchmarks: Hilti SLAM Challenge 2022/2023 and Newer College Dataset provide construction, campus, and underground sequences.
@@
Baselines: FAST-LIVO2, FAST-LIVO, R3LIVE (LiDAR-inertial-visual), FAST-LIO2 (LiDAR-inertial), and OpenMAVIS (multi-camera visual-inertial). All LIVO systems use 10Hz for both camera and LiDAR data acquisition. FAST-LIVO2 and our system use identical LIO and IMU parameters.
Omni-LIVO Variants: We evaluate five variants: w/o ac (Adaptive Covariance disabled), w/o cvtm (Cross-View Temporal Migration disabled), 1-Cam, 2-Cam, and Ours (complete system with three cameras).
IV-B Quantitative Results
Table II summarizes absolute trajectory errors (RMSE) across all sequences. Omni-LIVO variants achieve the lowest error in 25 out of 27 sequences and successfully complete all evaluated tracks, outperforming FAST-LIVO2, R3LIVE, FAST-LIO2, and OpenMAVIS across diverse environments including construction sites, indoor corridors, and underground scenarios.
@@ IV-B Quantitative Results
Ablation analysis demonstrates that Cross-View Temporal Migration improves accuracy in sequences with large viewpoint changes (Cupola 2: 0.098 m vs. 0.146 m w/o cvtm), while Adaptive Covariance enhances robustness under illumination variations (Lower Gallery 2: 0.083 m vs. 0.109 m w/o ac). Camera number comparison shows progressive gains from 1-Cam to 3-Cam (Ours), with the largest improvements in complex environments (Underground hard: 0.096 m to 0.055 m).
@@ IV-C Computational Performance
Performance evaluation is conducted on NVIDIA Jetson AGX Orin with 12-core ARM CPU (2.2GHz) and 64GB RAM running Ubuntu 20.04. System parameters: LiDAR voxel size 1.0 m, surface filtering threshold 0.5 m, maximum 150 visual points per frame.
@@
TABLE III: Computational Performance Comparison
Sequence #Cam FAST-LIVO2 Omni-LIVO
Time(ms) Mem(MB) Time(ms) Mem(MB)
Construction Upper 3 50.7 1901 58.1 2017
Cupola 2 3 51.2 3477 57.0 3498
Outside Building 3 54.5 3261 65.9 3203
Average (Hilti) 3 52.1 2880 60.3 2906
Corridor 4 17.3 1054 21.0 1197
Basement3 4 49.9 3473 60.1 3520
Average (Custom) 4 33.6 2264 40.6 2359
@@ V Conclusion
This letter presents Omni-LIVO, a tightly coupled multi-camera LiDAR-inertial-visual odometry system that extends direct visual methods to multi-view configurations through Cross-View temporal migration and adaptive ESIKF integration. Evaluation on public benchmarks and custom datasets shows that the system achieves lower trajectory errors compared to state-of-the-art LIVO, LIO, and visual-inertial methods. The multi-camera configuration provides extended spatial coverage and produces denser RGB-colored point clouds. Future work will explore online loop closure and investigate multi-sensor calibration refinement.

### v3

@@ Abstract
Wide field-of-view (FoV) LiDAR sensors provide dense geometry across large environments, but existing LiDAR-inertial-visual odometry (LIVO) systems generally rely on a single camera, limiting their ability to fully exploit LiDAR-derived depth for photometric alignment and scene colorization. We present Omni-LIVO, a tightly coupled multi-camera LIVO system that leverages multi-view observations to comprehensively utilize LiDAR geometric information across extended spatial regions. Omni-LIVO introduces a Cross-View direct alignment strategy that maintains photometric consistency across non-overlapping views, and extends the Error-State Iterated Kalman Filter (ESIKF) with multi-view updates and adaptive covariance. The system is evaluated on public benchmarks and our custom dataset, showing improved accuracy and robustness over state-of-the-art LIVO, LIO, and visual-inertial SLAM baselines.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) has become essential for autonomous systems operating in unknown environments. Traditional methods are categorized by sensor types: vision-based SLAM [1, 2] provides rich visual information but struggles in low-light or texture-poor scenes, while LiDAR-based SLAM [3] offers robust geometry but degrades in repetitive environments. IMUs, though environment-independent, suffer from long-term drift.
@@
Realizing these advantages requires addressing key technical challenges. Unlike stereo systems that use overlapping views for feature matching, non-overlapping multi-camera configurations require different approaches. First, photometric continuity must be maintained through temporal patch migration as patches transition between cameras. Second, adaptive covariance modeling is needed to handle heterogeneous measurement reliability across different views [11, 12]. Directly leveraging LiDAR-derived depth for multi-camera sparse direct alignment maintains efficient performance while enabling accurate visual point selection and extended spatial coverage beyond single-camera constraints.
To address these challenges, we present Omni-LIVO [^1] Code and data are available at https://github.com/elon876/Omni-LIVO., a tightly coupled multi-camera LIVO system that extends FAST-LIVO2 with multi-view photometric constraints.
@@
    Extensive experimental evaluation: Comprehensive evaluation of the tightly coupled multi-camera LIVO framework on public benchmarks and custom datasets.
@@ II-C Multi-Camera Visual-Inertial Systems
Multi-camera visual-inertial SLAM has gained attention for extended spatial coverage. MIMC-VINS [18] supports arbitrary camera numbers using multi-state constraint Kalman filtering. BAMF-SLAM [19] integrates multiple fisheye cameras with IMU through bundle adjustment. For large FoV coverage, LF-VIO [20] and LF-VISLAM [21] employ negative imaging plane models for single wide-FoV cameras. MAVIS [10] extends to partially overlapped multi-camera systems. These systems operate in the visual-inertial domain without leveraging LiDAR geometric structure.
@@ II-D Direct Visual Methods
Direct visual methods minimize photometric error without feature extraction. DTAM [22], LSD-SLAM [1], and DSO [23] established foundations for dense and sparse tracking. MultiCol-SLAM [24] extended direct methods to multi-camera systems. This suits multi-camera LIVO by avoiding correspondence complexity in non-overlapping views.
@@ III System Overview And Methodology
The Omni-LIVO system comprises four modules: Sensor Acquisition, Preprocessing, Multi-Camera ESIKF, and Unified Mapping, as illustrated in Fig. 1. The processing pipeline begins with IMU propagation, followed by sequential ESIKF updates that integrate LiDAR point-to-plane residuals with multi-view photometric constraints. Cross-View temporal migration and adaptive covariance preserve photometric consistency across non-overlapping views.
@@ III-B Multi-Sensor Data Alignment
Building upon FAST-LIVO2 [8], we extend temporal alignment to multi-camera configurations [25]. The system forms temporally coherent measurement packets:
@@ III-C Cross-View Direct Method
$\mathbf{r}_{VIO}=\left[\mathbf{r}^{1}_{\mathrm{direct}},\ldots,\mathbf{r}^{N_{c}}_{\mathrm{direct}},\mathbf{r}^{1\to 2}_{\mathrm{migration}},\ldots,\mathbf{r}^{i\to j}_{\mathrm{migration}}\right]^{T}$ (13)
@@ III-E Iterative Multi-Camera State Estimation with ESIKF
Measurement quality varies across cameras due to heterogeneous exposure and environments. We extend ESIKF to integrate multi-view photometric constraints with LiDAR measurements through sequential update and adaptive covariance, as shown in Fig. 5. The state is first updated with LiDAR point-to-plane residuals, then sequentially updated with multi-camera photometric residuals. Multi-camera residuals and Jacobians are stacked across all cameras and processed through iterative ESIKF updates until convergence.
@@
where $\bar{e}^{(t-1)}$ is the mean squared photometric error from the previous frame, $\sigma_{\min}$ and $\sigma_{\max}$ are the minimum and maximum covariance scaling bounds (set to $100.0$ and $1000.0$ respectively), and $e_{\max}$ is the maximum error threshold for scaling (typically set to $100.0$). Temporal smoothing ensures stability:
@@
Algorithm 1 summarizes the complete implementation. Algorithm-specific notation: $\check{\mathbf{x}}^{(\ell)}$ is the linearization point; $\mathbf{r}_{c}^{(\ell)}$, $\mathbf{H}_{c}^{(\ell)}$ are per-camera residuals and Jacobians stacked into $\mathbf{r}^{(\ell)}_{VIO}$, $\mathbf{H}^{(\ell)}_{VIO}$; $\delta\hat{\mathbf{x}}^{(\ell)}$ is the state increment; $e^{(\ell)}$ is the mean squared error; $\ell_{\max}=10$ is the maximum iteration count.
1IMU propagation: $\hat{\mathbf{x}}^{-}$, $\mathbf{P}^{-}$; LiDAR plane extraction
2 LiDAR update: Iterative ESIKF with $\mathbf{r}_{LIO}$, $\mathbf{H}_{LIO}$
3 Select patches from cameras $c\in[1,N_{c}]$; Compute $\tilde{\sigma}^{(t)}$ via Eq. (17)-(18)
4 VIO update: $\ell\leftarrow 0$, $\delta\hat{\mathbf{x}}^{(0)}\leftarrow\mathbf{0}$, $e_{\text{last}}\leftarrow\infty$
5 repeat
6    $\ell\leftarrow\ell+1$; Update $\check{\mathbf{x}}^{(\ell)}=\hat{\mathbf{x}}^{-}\boxplus\delta\hat{\mathbf{x}}^{(\ell-1)}$
7    for camera $c=1,\ldots,N_{c}$ do
8        Compute $\mathbf{r}_{c}^{(\ell)}$, $\mathbf{H}_{c}^{(\ell)}$ with cross-camera constraints
9
10   Stack $\mathbf{r}^{(\ell)}_{VIO}=[\mathbf{r}_{1}^{(\ell)};\ldots;\mathbf{r}_{N_{c}}^{(\ell)}]$, $\mathbf{H}^{(\ell)}_{VIO}=[\mathbf{H}_{1}^{(\ell)};\ldots;\mathbf{H}_{N_{c}}^{(\ell)}]$
11    Compute $e^{(\ell)}=\|\mathbf{r}^{(\ell)}_{VIO}\|^{2}/n_{\text{meas}}$
12    if $e^{(\ell)}\leq e_{\text{last}}$ then
13        $e_{\text{last}}\leftarrow e^{(\ell)}$
14        $\mathbf{K}=((\mathbf{H}^{(\ell)}_{VIO})^{\mathrm{T}}\mathbf{H}^{(\ell)}_{VIO}+\mathbf{P}^{-1}/\tilde{\sigma}^{(t)})^{-1}$
15        $\delta\hat{\mathbf{x}}^{(\ell)}=\mathbf{K}((\mathbf{H}^{(\ell)}_{VIO})^{\mathrm{T}}\mathbf{r}^{(\ell)}_{VIO}-\mathbf{P}^{-1}\delta\hat{\mathbf{x}}^{(\ell-1)}/\tilde{\sigma}^{(t)})$
16        $\hat{\mathbf{x}}^{-}\leftarrow\hat{\mathbf{x}}^{-}\boxplus\delta\hat{\mathbf{x}}^{(\ell)}$
17
18   else
19        break
20
21
22until $\|\delta\mathbf{R}^{(\ell)}\|<0.001^{\circ}$ and $\|\delta\mathbf{t}^{(\ell)}\|<0.01$$m$ or $\ell\geq\ell_{\max}$
[…]
return $\hat{\mathbf{x}}_{k|k}$, $\mathbf{P}_{k|k}$
@@ IV Experimental Evaluation
We evaluate the proposed system on public benchmarks and custom datasets, and compare with state-of-the-art methods.
@@ IV-A Datasets and Baselines
To demonstrate the effectiveness of the proposed multi-camera LIVO framework, we conduct comprehensive experiments evaluating trajectory accuracy, ablation of key components, computational efficiency, and qualitative mapping performance across diverse environments.
@@
We compare against five methods: FAST-LIVO2, FAST-LIVO, R3LIVE (LiDAR-inertial-visual), FAST-LIO2 (LiDAR-inertial), and OpenMAVIS (multi-camera visual-inertial). All LIVO systems use 10Hz for both camera and LiDAR data acquisition. FAST-LIVO2 and our system use identical LIO and IMU parameters. Ablation studies are conducted using five system variants: w/o ac (Adaptive Covariance disabled), w/o cvtm (Cross-View Temporal Migration disabled), 1-Cam, 2-Cam, and Ours (complete system with three cameras).
IV-B Quantitative Results
Table II summarizes absolute trajectory errors (bold: best, underline: second-best) across all sequences. Omni-LIVO variants achieve the lowest error in 25 out of 27 sequences and successfully complete all evaluated tracks, outperforming FAST-LIVO2, R3LIVE, FAST-LIO2, and OpenMAVIS across diverse environments including construction sites, indoor corridors, and underground scenarios. Compared to FAST-LIVO2, our system achieves an average 30% error reduction across all 27 sequences, with improvements up to 82% on Lower Gallery.
Ablation analysis demonstrates that Cross-View Temporal Migration improves accuracy in sequences with large viewpoint changes (Cupola 2: 0.098 m vs. 0.146 m w/o cvtm), while Adaptive Covariance enhances robustness under illumination variations (Lower Gallery 2: 0.083 m vs. 0.109 m w/o ac). Camera number comparison shows progressive gains from 1-Cam to 3-Cam (Ours), with the largest improvements in complex environments (Underground hard: 0.096 m to 0.055 m). This trend indicates that multi-camera fusion provides greater benefits in geometrically challenging scenarios.
@@ IV-C Computational Performance
We use NVIDIA Jetson AGX Orin with 12-core ARM CPU (2.2GHz) and 64GB RAM running Ubuntu 20.04. System parameters: LiDAR voxel size 1.0 m, surface filtering threshold 0.5 m, maximum 150 visual points per frame.
@@ IV-D Qualitative Results
TABLE III: Computational Performance Comparison
Sequence #Cam FAST-LIVO2 Omni-LIVO
Time Mem Time Mem
(ms) (MB) (ms) (MB)
Construction Upper 3 50.7 1901 58.1 2017
Cupola 2 3 51.2 3477 57.0 3498
Outside Building 3 54.5 3261 65.9 3203
Avg. (Hilti) 3 52.1 2880 60.3 2906
Corridor 4 17.3 1054 21.0 1197
Basement3 4 49.9 3473 60.1 3520
Avg. (Custom) 4 33.6 2264 40.6 2359
@@ V Conclusion
This letter presents Omni-LIVO, a tightly coupled multi-camera LiDAR-inertial-visual odometry system that extends direct visual methods to multi-view configurations through Cross-View temporal migration and adaptive ESIKF integration. Evaluation on public benchmarks and custom datasets shows that the system achieves lower trajectory errors compared to state-of-the-art LIVO, LIO, and visual-inertial methods. The multi-camera configuration provides extended spatial coverage and produces denser RGB-colored point clouds. The quantitative and qualitative improvements demonstrate the effectiveness of the proposed multi-camera fusion approach. Future work will explore online loop closure and investigate multi-sensor calibration refinement.